OpenStudy (anonymous):

Use DeMoivre’s Theorem to find [2(cosπ/10+ i sinπ/10)]^5in standard form. a) Identify the real part. b)identify the imaginary part.

OpenStudy (mertsj):

z=r(cos(x)+i sin(x)) then z^n=r^n(cos(nx)+isin(nx))

OpenStudy (mertsj):

So 32(cos( pi/2)+isin(pi/2))=32(0+i(1))=0+32i The real part is 0 and the imaginary part is 32i